How do you write the linear equation in slope-intercept form: 9x-7y=6
Answer by tutorcecilia(2144)


Use the formula for the slope-intercept equation of a line: y=mx+b
.
9x-7y=6 [solve for y or "isolate" the y-term]
9x-9x-7y=-9x+6
-7y=-9x+6
-7y/-7=-9x/-7+6/-7
y=9/7x-6/7